Two charges are on the x-axis. Charge q 1= +35 nC is at x = 3.00 cm while charge q 2= -65nC is at x = 6.00 cm. Find the magnitude and direction of


a) the electric force that q 1 exerts on q2 ; and


b) the electric force the q 2 exerts on q1 .


2. A positive charge of 10 -8 coulombs (C) experiences a force of 0.95 N when located at a certain point in an electric field. Find the intensity (magnitude) of the electric field at that point.


1
Expert's answer
2022-04-05T09:45:07-0400

(1)Force


"F_{12}=\frac{kq_1q_2}{r^2}=-\frac{9\times10^9\times35\times10^{-9}\times65\times10^{-9}}{0.03^2}=-0.02275N"

(2)Force

"F_{21}=\frac{kq_1q_2}{r^2}=-\frac{9\times10^9\times35\times10^{-9}\times65\times10^{-9}}{0.03^2}=-0.02275N"

2

Electric field

"E=\frac{F}{q}"

"E=\frac{0.95}{10^{-8}}=95\times10^6N/c"

Magnitude of electric field

"|E|=95\times10^6N/c"